# Distributed Maximum Flow in Planar Graphs

### v3

@@ 6.1 Maximum st-Flow
Here we desire to simulate [42]’s approximate SSSP algorithm on $(G\cup\{e\})^{*}\neq G^{*}$ by applying Section 4 (Theorem B.14) to simulate the extended minor aggregation model on $G^{*}$. We exploit that the extended model can simulate any basic or extended minor aggregation algorithm on any virtual graph $G^{*}_{virt}$ that contains a small (poly-logarithmic) number of arbitrarily connected virtual nodes (see Appendix B for exact definitions). Thus, the only thing we need to do, is to use that (Theorem B.14) in order to apply the approximate SSSP algorithm on $G^{*}_{virt}=(G\cup\{e\})^{*}$. To do so, we only need to distributively store $(G\cup\{e\})^{*}$ as needed (matching the input format of the algorithm stated in Theorem B.14, for full details, see Appendix B). We explain; Recall, the face-disjoint graph $\hat{G}$ is used as the base graph for representing $G^{*}$, i.e., $\hat{G}$ is the distributed method for storing $G^{*}$; To store any virtual graph $G^{*}_{virt}$ that is obtained by replacing nodes or adding virtual nodes to $G^{*}$, we need that all vertices on a face $\hat{g}$ of $\hat{G}$ (maps to node $g\in G^{*}$) to know a list of (IDs of) $g$’s incident virtual edges and all virtual edges that connect any two virtual nodes (and their weight if exists) in $G^{*}_{virt}$. So, we need to (1) show how to delete $f$ from $G^{*}$, (2) add two nodes that play the roles of $f_{1},f_{2}$, and (3), connect them to neighbors of $f$ and to themselves properly, represented in $\hat{G}$ as described; Such that, the obtained $G^{*}_{virt}$ has the exact same structure as $(G\cup\{e\})^{*}$. Note, changes need only to be made on $\hat{f}$’s vertices and vertices lying on (faces $\hat{g}$ in $\hat{G}$ corresponding to) neighbors $g$ of $f$ in $G^{*}$.
@@
We start by: (1) Enumerating $\hat{f}$’s vertices in $\hat{G}$ in ascending order and deactivating them (deleting $f$). This numbering would be useful to define $f_{1},f_{2}$ (i.e, split $\hat{f}$). (2) Learning (locally) in $\hat{G}$ for each edge of $E_{C}$  [^8] Those edges map 1-1 to edges of $G^{*}$, see Section 3 and LABEL:appendix:_preliminaries_hat{G} for more details. that is incident to $\hat{f}$, whether its corresponding edge in $G^{*}$ shall be incident to $f_{1}$ or to $f_{2}$. Then, (3) Each (face $\hat{g}$ in $\hat{G}$ that maps to a) neighbor $g$ of $f$ needs to know its minimal weight edge connecting it to $f_{1}$ and $f_{2}$, as we are interested in shortest paths and cannot afford to broadcast all edges of $f$, due to their possibly large quantity. To do so, we compute a minimum aggregate over weights of $\hat{g}$’s edges that are mapped to edges that shall be incident to $f_{1}$ ($f_{2}$), so that all vertices in $\hat{g}$ know the edge that connects $g$ to $f_{1}$ ($f_{2}$). Finally, (4) A global maximum operator is computed over edge weights of $G$, such that, all vertices of $\hat{G}$ learn that $f_{1},f_{2}$ are connected by an edge whose weight is $nW$, where $W$ is the maximum weight found. We describe how each step is implemented; Assume $\hat{G}$ is already constructed and its faces that correspond to $G$ faces are identified (Properties 1, 4 of $\hat{G}$).
@@
    For enumerating $\hat{f}$’s vertices in $\hat{G}$, we use a simple classic procedure of a subtree sum using part-wise aggregations on $\hat{G}$. Specifically, the input tree is $\hat{f}$ minus the $E_{R}$  [^9] For each edge $e$ of $G$, there are two edge copies in $E_{R}\subset E(\hat{G})$. Intuitively, there is an edge copy for each face of $G$ that contains $e$. edge copy of the edge preceding $e$ in $s$’s local ordering of its incident edges (this edge is known to $s$ locally), the root is the copy of $s$ that is incident to the removed edge, and all vertices get an input value of one initially. The procedure we use is similar to those described in [14, 17]. We can do so in $\tilde{O}(D)$ rounds by Section 4 (Lemma B.6).
    Denote by $s^{\prime}$, the copy of $s$ in $\hat{f}$ which is incident to the two consecutive edges $e_{1}^{\prime},e_{2}^{\prime}$ (of $E_{R}$ in $\hat{G}$), that map to the edges $e_{1},e_{2}$ incident to $s$, s.t. $e$ is embedded in between $e_{1},e_{2}$ (in $G$), we define $t^{\prime}$ analogously. Intuitively, let $e_{a},e_{b}$ be the two edges incident to $s$ s.t. all edges between $e_{a},e_{b}$ (in the local ordering of $s$) are in $f$, then all edges in between $e_{a},e_{1}$ have to go to $f_{1}$ and all edges in between $e_{2},e_{b}$ have to go to $f_{2}$, as $e$ is embedded right after $e_{1}$ and right before $e_{2}$. We define $s^{\prime}$ to capture this in $\hat{G}$ (where edges have a lot of copies, sometimes in the same face  [^10] E.g, if an edge $e$ is internal to a face $f$, then it has its two $E_{R}$ copies in the face $\hat{f}$ of $\hat{G}$ that maps to $f$. More generally, if there is a tree attatched to a vertex in the interior of a face $f$, then each of those tree edges has two copies in $\hat{f}$, as $\hat{f}$ is basically an Euler tour on $f$. See Section 3 and LABEL:appendix:_preliminaries_hat{G} for more details.), in order to define $f_{1},f_{2}$ correctly.
@@
Notice, all the above terminates within $\tilde{O}(D)$ rounds on $G$ by Section 4 (Lemma B.6). Finally, since $G^{*}_{virt}=(G\cup\{e\})^{*}$ now is properly distributively stored, then, we can simulate the extended minor-aggregation model on it (see Theorem B.14 for details); So, we first use it in order to deactivate parallel edges (Lemma B.15), keeping only the minimal-weighted edge between its endpoints (we aim to compute SSSP). Then, since $G^{*}_{virt}$ is simple and we can simulate any minor aggregation algorithm on it, it satisfies the requirements for applying the mentioned theorem with [42]’s SSSP algorithm, obtaining an approximate SSSP tree rooted at the source $f_{1}$.
@@
Assuming we have computed an approximate SSSP tree $T$ (with zero weighted edges included) as promised, we can compute the distances of any node from the source by a simple procedure. That is, root the computed tree with the source, then compute the tree ancestor sum of every node $g\in T$, i.e., the weight of the $f_{1}$-to-$g$ $T$-path. For that, we use a procedure of [18] that runs in $\tilde{O}(1)$ minor aggregation rounds (Lemma 16 in their paper), thus, $\tilde{O}(D)$ rounds in $G$. After that process terminates, each vertex lying on $f_{2}$ (in $G$) knows the approximate distance from $f_{1}$ to $f_{2}$ in $G^{*}_{virt}$, thus, it can broadcast it to the whole graph. Note, that distance is a $(1+\epsilon)$ approximation of the maximum $st$-flow value (by Hassin [20]). Since the algorithm guarantees to report $\alpha\cdot\mathsf{dist}(f_{1},f_{2})$ instead of $\mathsf{dist}(f_{1},f_{2})$, for $\alpha\in[1,1+\epsilon]$, multiplying by $(1-\epsilon)$, we get $(1-\epsilon)\alpha\in[1-\epsilon,1-\epsilon^{2}]$, that is, a $(1-\epsilon)$ approximation of the maximum $st$-flow value.
@@
Note that in Theorem 1.2 we obtained both the value of the maximum flow and a corresponding assignment of flow to the edges. Here however, we approximate the flow value but do not immediately obtain an assignment (or an approximate assignment) of the flow. The reason is as follows: Hassin’s algorithm was designed as an exact algorithm (running an exact SSSP computation) while we use an approximate SSSP. For the flow value this is fine, instead of $\mathsf{dist}(f_{1},f_{2})$ we report a $(1-\epsilon)$-approximation of $\mathsf{dist}(f_{1},f_{2})$. However, for the corresponding flow assignment, recall that the flow on edge $e^{\prime}\in G$ with $e^{\prime*}=(f,g)\in(G\cup\{e\})^{*}$ is $\mathsf{dist}(f_{1},g)-\mathsf{dist}(f_{1},f)$. This means we need the approximation algorithm to bound $\mathsf{dist}(f_{1},g)-\mathsf{dist}(f_{1},f)$ by the edge’s capacity (and not just approximate $\mathsf{dist}(f_{1},g)$ and $\mathsf{dist}(f_{1},f)$). I.e., we want the approximate distances to satisfy the triangle inequality. This indeed is possible using (a dual implementation of) a recent algorithm of Rozhon, Haeupler, Martinsson, Grunau and Zuzic [40], however, their algorithm was not implemented in the minor-aggregation model. Given access to a $(1+\epsilon)$-approximate shortest paths oracle, [40]’s algorithm produces smooth (defined next) approximate distances within a logarithmic number of calls to the oracle with a parameter $\epsilon^{\prime}=O(\epsilon/\log n)$. Before we dive into the simulation, we define what a smooth distance approximation is, suggest a flow assignment and prove it is a feasible $(1-\epsilon)$ approximate assignment using the smoothness property.
@@
    Conservation. By [25], assuming a planar directed graph with edge capacities and an assignment of potentials (weights) $\rho(\cdot)$ on its faces, then, assigning for each edge a flow of $\rho(g)-\rho(f)$, where $g$ (resp. $f$) is the face on the right (left) of that edge, gives a flow assignment that obeys conservation in all vertices  [^11] This kind of a flow assignment is referred to as a circulation in the centralized literature on planar graphs.. Considering $G\cup\{e\}$, the assignment we suggest, is deliberately defined so that the directions of the flow would be consistent with the above. I.e, directing an edge in its defined flow direction and pushing $|\delta(g)-\delta(f)|$ units of flow would be consistent with the definition above. Thus, the flow assignment we suggest obeys conservation in every vertex of $G\cup\{e\}$. Note, $G$ is equivalent to $G\cup\{e\}$ after removing $e=\{s,t\}$, thus, the induced flow assignment on $G$ obeys conservation except in the endpoints of $e$, which are $s$ and $t$.
@@
that we can apply [40]’s algorithm (Theorem 5.1, Algorithms 3 and 4 in their paper), in order to obtain smooth $(1+\epsilon)$ approximate single source shortest paths from $f_{1}$ in $G^{*}_{virt}=(G\cup\{e\})^{*}$ by invoking the algorithm of [42] that we use for approximating distances, which we refer to as the oracle. We note that [40]’s algorithm assumes real weights in $[1,n^{O(1)}]$, whereas we work with integral weights, however, they suggest a fix (by scaling) for this that appears in Appendix C of their paper. Note, scaling can be implemented efficiently in one minor-aggregation round, henceforth, we consider this issue resolved.
@@
[40]’s algorithm in a high-level. We describe [40]’s algorithm for smooth distance approximation in a high-level, highlighting only the computational tasks necessary to implement it in the minor aggregation model. For the algorithm in full resolution, the reader is referred to Section 5 of [40]. Initially, nodes of the input graph shall know the source $s$, the approximation parameter $\epsilon$ and a trivial estimate $\Delta$ of the maximum distance in the graph ($\Delta:=2nW$ in our case), then, each node computes additional parameters that depend only on the former. Note, all of the former may be implemented in only one minor-aggregation round.
@@
Next, the algorithm computes a distance approximation, based on a call to the approximation oracle (any SSSP approximation algorithm, without any additional special guarantees) with $s$ and $\epsilon^{\prime}=O(\epsilon/\log n)$ Then, sequentially, $O(\log n)$ phases are performed, each defining a distance approximation, such that, the final output of the algorithm is the approximation computed in the last phase. In each phase, two calls to the approximate oracle are performed with $s$ and $\epsilon^{\prime}/100$. Each call to the oracle is on a level graph. Afterwards, each level graph defines a distance approximation of the input graph, and the phase’s distance approximation is computed by taking the element- (node-) wise minimum over: (1) the approximations computed on the level graphs, and (2) the approximation from the previous phase. Again, taking the minimum requires only one minor-aggregation round.
@@
Level graphs. The only vagueness in the implementation lies in the level graphs, which we now define, then, show that they can constructed and simulated efficiently in the minor-aggregation model. The level graph $H$ (Definition 5.7 in [40]) in a specific phase of an input graph $\mathcal{G}$ where each node of $\mathcal{G}$ knows its (approximate) distance from a source $s$, is defined as follows. The node and edge sets of $H$ are the same as those of $\mathcal{G}$, then: (1) Each node assigns itself a level (a number), depending only on its distance from $s$ and some parameter $\omega$, (2) each edge that connects neighbors of different levels is removed, the weights of edges kept are altered depending only on previously computed parameters that all nodes of $\mathcal{G}$ know. Finally, (3) An edge between $s$ and each $u\in\mathcal{G}$ is added, with weight depending only on $\omega$ and the $s$-to-$u$ distance approximated in the previous phase.
@@
Note, each phase is of $\tilde{O}(1)+T_{SSSP(\epsilon^{\prime})}$ minor-aggregation rounds, where $T_{SSSP(\epsilon^{\prime})}$ denotes the minor-aggregation round complexity of a $(1+\epsilon^{\prime})$-approximate SSSP oracle. Thus, the overall minor-aggregate round complexity of the $(1+\epsilon)$ smooth approximate distance computation is $O(\log n)\cdot\tilde{O}(T_{SSSP(\epsilon^{\prime})})$. Implementing the approximate SSSP oracle with the algorithm of [42], results with a total minor-aggregate round complexity of $O(\log n)\cdot(\frac{1}{\epsilon^{\prime}})^{2}\cdot 2^{O(\log n\cdot\log\log n)^{3/4}}$, i.e., $\frac{\log^{3}n}{\epsilon^{2}}\cdot 2^{O(\log n\cdot\log\log n)^{3/4}}$, which translates to $\frac{D}{\epsilon^{2}}\cdot n^{o(1)}$ $\mathsf{CONGEST}$ rounds on the network of communication $G$. For conclusion, since each vertex $v$ in $G$ simulates its copies in $\hat{G}$, which in turn simulate the faces $v$ participates in in $G^{*}_{virt}$, then, $v$ knows for each face it participates in, its distance from the source. Thus, after computing smooth distances in the relevant graph $G^{*}_{virt}$, in one round of communication, $v$ knows for each incident edge $e^{\prime}$, such that $e^{\prime*}=(f,g)$, the values $\delta(f_{1},g)$ and $\delta(f_{1},f)$ and can locally compute $\phi(e^{\prime})$. Finally, we are ready to go through the last modification required to keep our promise, that is, we show next how to allow zero capacities (zero-weight edges in the dual shortest paths approximation).
@@
The algorithm of [42] that we use for the approximate SSSP and of [40] used to obtain the smoothness property work for weights in $[1,n^{O(1)}]$, indeed all weights in $G^{*}_{virt}$, even that of the infinite capacity edge are in $[1,n^{O(1)}]$ except the zero-weighted ones. To deal with that issue, we simulate the obtained smooth approximate SSSP algorithm above while considering zero weighted edges as follows: We run the algorithm on maximal connected components of $G^{*}_{virt}\setminus\{e:w(e)=0\}$, obtaining a partial SSSP tree of $G^{*}_{virt}$, then we augment it with some zero weighted edges without closing a cycle, obtaining an SSSP tree $T$ of $G^{*}_{virt}$, all done via the extended minor aggregation model.
@@
Note, $T$ obeys the triangle-inequality (which is crucial for a flow assignment) for the following reasoning. Distances in $T_{1}$ represent distances in $G^{*}_{virt}$, by construction any pair of nodes in $T_{1}$ satisfies the triangle inequality. Thus, map each node $f\in G^{*}_{virt}$ to the super-node $a$ of $T_{1}$ that contains it, the property follows since $f$ and $a$ have the same distances from the source because $T_{1}$ was obtained by contracting zero-weight edges only. I.e, $T=T_{1}\cup T_{2}$ satisfies the triangle inequality as any pair of (super-)nodes in $T_{1}$ does. ∎

### v4

@@ 6.1 Maximum st-Flow
Here we desire to simulate [43]’s approximate SSSP algorithm on $(G\cup\{e\})^{*}\neq G^{*}$ by applying Section 4 (Theorem 4.14) to simulate the extended minor aggregation model on $G^{*}$. We exploit that the extended model can simulate any basic or extended minor aggregation algorithm on any virtual graph $G^{*}_{virt}$ that contains a small (poly-logarithmic) number of arbitrarily connected virtual nodes (see Section 4 for exact definitions). Thus, the only thing we need to do, is to use that (Theorem 4.14) in order to apply the approximate SSSP algorithm on $G^{*}_{virt}=(G\cup\{e\})^{*}$. To do so, we only need to distributively store $(G\cup\{e\})^{*}$ as needed (matching the input format of the algorithm stated in Theorem 4.14). We explain; Recall, the face-disjoint graph $\hat{G}$ is used as the base graph for representing $G^{*}$, i.e., $\hat{G}$ is the distributed method for storing $G^{*}$; To store any virtual graph $G^{*}_{virt}$ that is obtained by replacing nodes or adding virtual nodes to $G^{*}$, we need that all vertices on a face $\hat{g}$ of $\hat{G}$ (maps to node $g\in G^{*}$) to know a list of (IDs of) $g$’s incident virtual edges and all virtual edges that connect any two virtual nodes (and their weight if exists) in $G^{*}_{virt}$. So, we need to (1) show how to delete $f$ from $G^{*}$, (2) add two nodes that play the roles of $f_{1},f_{2}$, and (3), connect them to neighbors of $f$ and to themselves properly, represented in $\hat{G}$ as described; Such that, the obtained $G^{*}_{virt}$ has the exact same structure as $(G\cup\{e\})^{*}$. Note, changes need only to be made on $\hat{f}$’s vertices and vertices lying on (faces $\hat{g}$ in $\hat{G}$ corresponding to) neighbors $g$ of $f$ in $G^{*}$.
@@
We start by: (1) Enumerating $\hat{f}$’s vertices in $\hat{G}$ in ascending order and deactivating them (deleting $f$). This enumeration would be useful to define $f_{1},f_{2}$ (i.e, split $\hat{f}$). (2) Learning (locally) in $\hat{G}$ for each edge of $E_{C}$  [^14] Those edges map 1-1 to edges of $G^{*}$, see Section 3 and LABEL:appendix:_preliminaries_hat{G} for more details. that is incident to $\hat{f}$, whether its corresponding edge in $G^{*}$ shall be incident to $f_{1}$ or to $f_{2}$. Then, (3) Each (face $\hat{g}$ in $\hat{G}$ that maps to a) neighbor $g$ of $f$ needs to know its minimal weight edge connecting it to $f_{1}$ and $f_{2}$, as we are interested in shortest paths and cannot afford to broadcast all edges of $f$, due to their possibly large quantity. To do so, we compute a minimum aggregate over weights of $\hat{g}$’s edges that are mapped to edges that shall be incident to $f_{1}$ ($f_{2}$), so that all vertices in $\hat{g}$ know the edge that connects $g$ to $f_{1}$ ($f_{2}$). Finally, (4) A global maximum operator is computed over edge weights of $G$, such that, all vertices of $\hat{G}$ learn that $f_{1},f_{2}$ are connected by an edge whose weight is $nW$, where $W$ is the maximum weight found. We describe how each step is implemented; Assume $\hat{G}$ is already constructed and its faces that correspond to $G$ faces are identified (Properties 1, 4 of $\hat{G}$).
@@
[…]
    Denote by $s^{\prime}$, the copy of $s$ in $\hat{f}$ which is incident to the two consecutive edges $e_{1}^{\prime},e_{2}^{\prime}$ (of $E_{R}$ in $\hat{G}$), that map to the edges $e_{1},e_{2}$ incident to $s$, s.t. $e=(t,s)$ is embedded in between $e_{1},e_{2}$ (in $G$), we define $t^{\prime}$ analogously. Let $e_{a},e_{b}$ be the two edges incident to $s$ s.t. all edges between $e_{a},e_{b}$ (in the local ordering of $s$) are in $f$, then all edges in between $e_{a},e_{1}$ have to go to $f_{1}$ and all edges in between $e_{2},e_{b}$ have to go to $f_{2}$, as $e$ is embedded right after $e_{1}$ and right before $e_{2}$. We define $s^{\prime}$ to capture this in $\hat{G}$ (where edges have a lot of copies, sometimes in the same face  [^16] E.g, if an edge $e$ is internal to a face $f$, then it has its two $E_{R}$ copies in the face $\hat{f}$ of $\hat{G}$ that maps to $f$. More generally, if there is a tree attatched to a vertex in the interior of a face $f$, then each of those tree edges has two copies in $\hat{f}$, as $\hat{f}$ is basically an Euler tour on $f$. See Section 3 and LABEL:appendix:_preliminaries_hat{G} for more details.), in order to define $f_{1},f_{2}$ correctly.
@@
Notice, all the above terminates within $\tilde{O}(D)$ rounds on $G$ by Section 4 (Lemma 4.6). Finally, since $G^{*}_{virt}=(G\cup\{e\})^{*}$ now is properly distributively stored, then, we can simulate the extended minor-aggregation model on it (see Theorem 4.14 for details); So, we first use it in order to deactivate parallel edges (Lemma 4.15), keeping only the minimal-weighted edge between its endpoints (we aim to compute SSSP). Then, since $G^{*}_{virt}$ is simple and we can simulate any minor aggregation algorithm on it, it satisfies the requirements for applying the mentioned theorem with [43]’s SSSP algorithm, obtaining an approximate SSSP tree rooted at the source $f_{1}$.
@@
Assuming we have computed an approximate SSSP tree $T$ (with zero weighted edges included) as promised, we can compute the distances of any node from the source by a simple procedure. That is, root the computed tree with the source, then compute the tree ancestor sum of every node $g\in T$, i.e., the weight of the $f_{1}$-to-$g$ $T$-path. For that, we use a procedure of [18] that runs in $\tilde{O}(1)$ minor aggregation rounds (Lemma 16 in their paper), thus, $\tilde{O}(D)$ rounds in $G$ (Theorem 4.14). After that process terminates, each vertex lying on $f_{2}$ (in $G$) knows the approximate distance from $f_{1}$ to $f_{2}$ in $G^{*}_{virt}$, thus, it can broadcast it to the whole graph. Note, that distance is a $(1+\epsilon)$ approximation of the maximum $st$-flow value (by Hassin [20]). Since the algorithm guarantees to report $\alpha\cdot\mathsf{dist}(f_{1},f_{2})$ instead of $\mathsf{dist}(f_{1},f_{2})$, for $\alpha\in[1,1+\epsilon]$, multiplying by $(1-\epsilon)$, we get $(1-\epsilon)\alpha\in[1-\epsilon,1-\epsilon^{2}]$, that is, a $(1-\epsilon)$ approximation of the maximum $st$-flow value.
@@
Note that in Theorem 1.2 we obtained both the value of the maximum flow and a corresponding assignment of flow to the edges. Here however, we approximate the flow value but do not immediately obtain an assignment (or an approximate assignment) of the flow. The reason is as follows: Hassin’s algorithm was designed as an exact algorithm (running an exact SSSP computation) while we use an approximate SSSP. For the flow value this is fine, instead of $\mathsf{dist}(f_{1},f_{2})$ we report a $(1-\epsilon)$-approximation of $\mathsf{dist}(f_{1},f_{2})$. However, for the corresponding flow assignment, recall that the flow on edge $e^{\prime}\in G$ with $e^{\prime*}=(g,h)\in(G\cup\{e\})^{*}$ is $\mathsf{dist}(f_{1},h)-\mathsf{dist}(f_{1},g)$. This means we need the approximation algorithm to bound $\mathsf{dist}(f_{1},h)-\mathsf{dist}(f_{1},g)$ by the edge’s capacity (and not just approximate $\mathsf{dist}(f_{1},h)$ and $\mathsf{dist}(f_{1},g)$). I.e., we want the approximate distances to satisfy the triangle inequality. This indeed is possible using (a dual implementation of) a recent algorithm of Rozhon, Haeupler, Martinsson, Grunau and Zuzic [41], however, their algorithm was not implemented in the minor-aggregation model. Given access to a $(1+\epsilon)$-approximate shortest paths oracle, [41]’s algorithm produces smooth (defined next) approximate distances within a logarithmic number of calls to the oracle with a parameter $\epsilon^{\prime}=O(\epsilon/\log n)$. Before we dive into the simulation, we define what a smooth distance approximation is, suggest a flow assignment and prove it is a feasible $(1-\epsilon)$ approximate assignment using the smoothness property.
@@
    Conservation. By [26], assuming a directed planar graph with edge capacities and an assignment of potentials (weights) $\rho(\cdot)$ on its faces, then, assigning for each edge a flow of $\rho(g)-\rho(f)$, where $g$ (resp. $f$) is the face on the right (left) of that edge, gives a flow assignment that obeys conservation in all vertices  [^17] This kind of a flow assignment is referred to as a circulation in the centralized literature on planar graphs.. Considering $G\cup\{e\}$, the assignment we suggest, is deliberately defined so that the directions of the flow would be consistent with the above. I.e, directing an edge in its defined flow direction and pushing $|\delta(g)-\delta(f)|$ units of flow would be consistent with the definition above. Thus, the flow assignment we suggest obeys conservation in every vertex of $G\cup\{e\}$. Note, $G$ is equivalent to $G\cup\{e\}$ after removing $e=\{s,t\}$, thus, the induced flow assignment on $G$ obeys conservation except in the endpoints of $e$, which are $s$ and $t$.
@@
It remains only to show that we can apply [41]’s algorithm (Theorem 5.1, Algorithms 3 and 4 in their paper), in order to obtain smooth $(1+\epsilon)$ approximate single source shortest paths from $f_{1}$ in $G^{*}_{virt}=(G\cup\{e\})^{*}$ by invoking the algorithm of [43] that we use for approximating distances, which we refer to as the oracle. We note that [41]’s algorithm assumes real weights in $1,n^{O(1)}]$, whereas we work with integral weights, however, they suggest a fix (by scaling) for this problem that appears in Appendix C of their paper. Note, scaling can be implemented efficiently in one minor-aggregation round, henceforth, we consider this issue resolved.
@@
[41]’s algorithm in a high-level. We describe [41]’s algorithm for smooth distance approximation in a high-level, highlighting only the computational tasks necessary to implement it in the minor aggregation model. For the algorithm in full resolution, the reader is referred to Section 5 of [41]. Initially, nodes of the input graph shall know the source $s$, the approximation parameter $\epsilon$ and a trivial estimate $\Delta$ of the maximum distance in the graph ($\Delta:=2nW$ in our case), then, each node computes additional parameters that depend only on the former. Note, all of the former is implemented in only one minor-aggregation round.
@@
Next, the algorithm computes a distance approximation, based on a call to the approximation oracle (any SSSP approximation algorithm, without any additional special guarantees) with $s$ and $\epsilon^{\prime}=O(\epsilon/\log n)$ . Then, sequentially, $O(\log n)$ phases are performed, each defining a distance approximation, such that, the final output of the algorithm is the approximation computed in the last phase. In each phase, two calls to the approximate oracle are performed with $s$ and $\epsilon^{\prime}/100$. Each call to the oracle is on a level graph. Afterwards, each level graph defines a distance approximation of the input graph, and the phase’s distance approximation is computed by taking the element- (node-) wise minimum over: (1) the approximations computed on the level graphs, and (2) the approximation from the previous phase. Again, taking the minimum requires only one minor-aggregation round.
@@
Level graphs. The only vagueness in the implementation lies in the level graphs, which we now define, then, show that they can be constructed and simulated efficiently in the minor-aggregation model. The level graph $H$ (Definition 5.7 in [41]) of an input graph $\mathcal{G}$ where each node of $\mathcal{G}$ knows its (approximate) distance from a source $s$, in a specific phase, is defined as follows. The node and edge sets of $H$ are the same as those of $\mathcal{G}$, then: (1) Each node assigns itself a level (a number), depending only on its distance from $s$ and some parameter $\omega$, (2) each edge that connects neighbors of different levels is removed, the weights of edges kept are altered depending only on previously computed parameters that all nodes of $\mathcal{G}$ know. Finally, (3) An edge between $s$ and each $u\in\mathcal{G}$ is added, with weight depending only on $\omega$ and the $s$-to-$u$ distance approximated in the previous phase.
@@
Note, each of the $O(\log n)$ phases is of $\tilde{O}(1)+T_{SSSP(\epsilon^{\prime})}$ minor-aggregation rounds, where $T_{SSSP(\epsilon^{\prime})}$ denotes the minor-aggregation round complexity of a $(1+\epsilon^{\prime})$-approximate SSSP oracle. Thus, the overall minor-aggregate round complexity of the $(1+\epsilon)$ smooth approximate distance computation is $O(\log n)\cdot\tilde{O}(T_{SSSP(\epsilon^{\prime})})$ for $\epsilon^{\prime}=(\epsilon/\log n)$. Implementing the approximate SSSP oracle with the algorithm of [43], results with a total minor-aggregate round complexity of $O(\log n)\cdot(\frac{1}{\epsilon^{\prime}})^{2}\cdot 2^{O(\log n\cdot\log\log n)^{3/4}}$, i.e., $\frac{\log^{3}n}{\epsilon^{2}}\cdot 2^{O(\log n\cdot\log\log n)^{3/4}}$, which by Theorem 4.14 translates to $\frac{D}{\epsilon^{2}}\cdot n^{o(1)}=D\cdot n^{o(1)}$ $\mathsf{CONGEST}$ rounds on the network of communication $G$, if we set $\epsilon=1/n^{o(1)}$. For conclusion, since each vertex $v$ in $G$ simulates its copies in $\hat{G}$, which in turn simulate the faces $v$ participates in in $G^{*}_{virt}$, then, $v$ knows for each face it participates in, its distance from the source. Thus, after computing smooth distances in the relevant graph $G^{*}_{virt}$, in one round of communication, $v$ knows for each incident edge $e^{\prime}$, such that $e^{\prime*}=(f,g)$, the values $\delta(f_{1},g)$ and $\delta(f_{1},f)$ and can locally compute $\phi(e^{\prime})$. Finally, we are ready to go through the last modification required to keep our promise, that is, we show next how to allow zero capacities (zero-weight edges in the dual shortest paths approximation).
@@
The algorithm of [43] that we use for the approximate SSSP and of [41] used to obtain the smoothness property work for weights in $1,n^{O(1)}]$, indeed all weights in $G^{*}_{virt}$, even that of the infinite capacity edge are in $1,n^{O(1)}]$ except the zero-weighted ones. To deal with that issue, we simulate the obtained smooth approximate SSSP algorithm above while considering zero weighted edges as follows: We run the algorithm on maximal connected components of $G^{*}_{virt}\setminus\{e:w(e)=0\}$, obtaining a partial SSSP tree of $G^{*}_{virt}$, then we augment it with some zero weighted edges without closing a cycle, obtaining an SSSP tree $T$ of $G^{*}_{virt}$, all done via the extended minor aggregation model.
@@
Note, $T$ obeys the triangle-inequality (which is crucial for a flow assignment) for the following reasoning. Distances in $T_{1}$ represent distances in $G^{*}_{virt}$, by construction any pair of nodes in $T_{1}$ satisfies the triangle inequality. Thus, map each node $f\in G^{*}_{virt}$ to the super-node $a$ of $T_{1}$ that contains it, the property follows since $f$ and $a$ have the same distances from the source because $T_{1}$ was obtained by contracting zero-weight edges only. I.e, $T=T_{1}\cup T_{2}$ satisfies the triangle inequality because each pair of (super-)nodes in $T_{1}$ does. ∎